xx TestTable – A Test Driver for the Key/Value Tables API

TestTable – A Test Driver for the Key/Value Tables API

Jonathan Thornburg <jthorn@aei.mpg.de>

Date

Abstract

This thorn is a test driver for the Util_Table*() key/valye-tables API.

1 Introduction

Cactus defines a set of Util_Table*() functions to manipulate key/value tables. (ChapterxC of the Cactus Users’ Guide contains a tutorial description of how to use these.)

This thorn is a test driver for these functions. It has no parameters; just activate it to run the tests. It will print out a bunch of messages (so you can see where it was if something goes wrong). If Cactus exits normally all the tests passed; any test failures will result in an assert() failure (which typically gives a core dump).

2 Parameters

3 Interfaces

General Implements: testtable

4 Schedule This section lists all the variables which are assigned storage by thorn CactusTest/TestTable. Storage can either last for the duration of the run (Always means that if this thorn is activated storage will be assigned, Conditional means that if this thorn is activated storage will be assigned for the duration of the run if some condition is met), or can be turned on for the duration of a schedule function.

Storage NONE

Scheduled Functions

CCTKxSTARTUP x testtablexdriver x test key/value tables api

11 x 11 Language: 11 c
11 x 11 Type:  11 function